OpenStudy (anonymous):

Name the property of real numbers illustrated by the equation. -4(x + 3) = -4x - 12 A. Commutative Property of Addition B. Associative Property of Multiplication C. Distributive Property D. Associative Property of Addition

OpenStudy (haseeb96):

A . is the correct answer because commutative property lie in this manner a + b = b + a and also order does not matter

OpenStudy (anonymous):

Thanks :)

OpenStudy (anonymous):

no a is not the correct answer

OpenStudy (anonymous):

it distributive due to distributing the -4 in the the perenthesis

OpenStudy (anonymous):

Oh...thanks :)

OpenStudy (haseeb96):

explain in more details

OpenStudy (anonymous):

well your multiplying not adding

OpenStudy (haseeb96):

distributive forms is A(B + C) = AB + AC

OpenStudy (anonymous):

and.....

OpenStudy (haseeb96):

yes u are right

OpenStudy (anonymous):

ya it is distributive

OpenStudy (anonymous):

lol its alright